OpenStudy (anonymous):

To reduce 16 ounces of a 25% solution of antiseptic to a 10% solution, how much distilled water should a nurse add? 12 oz 24 oz 36 oz

OpenStudy (anonymous):

Let X represent the water. 16 * 0.25 = (16+X)*.10 4 = 1.6 + .1X 2.4 = .1X 24 = X

OpenStudy (anonymous):

spanks